OpenStudy (anonymous):

@terenzreignz please help with this again.. Something different tho, confused on it. Not sure if you combine like terms or what. \[\sqrt[6]{27a^9b^6c^3}\] Simplify the expression.

OpenStudy (amistre64):

use the property:\[\Huge \sqrt[k]{a^n}=a^{(n/k)}\]

OpenStudy (anonymous):

Confused.. I get what k, a, and n equal, just not too sure of what to after the = sign of what you put... Lol. Sorry, new to this and kind'a not too good with it.. :/

OpenStudy (amistre64):

divide the exponents under the radical, by the number of the radical

OpenStudy (amistre64):

\[\Huge\sqrt[6]{27a^9b^6c^3}\] \[\Huge{27^{1/6}~a^{9/6}~b^{6/6}~c^{3/6}}\] then simplify those fractional exponents, preferably into mixed numerals

OpenStudy (anonymous):

Okay, so... \[27^{1/3}a1^{3/6}b^1c^{1/2}\] Or am I wrong? >.<

OpenStudy (amistre64):

youre on the right track, youve just got to get better skills with fractions :) 9/6 = 3/2 = 1 1/2 6/6 = 1 you got right 3/6 = 1/2 you got right 1/6 does not reduce to 1/3 tho \[\Huge{27^{1/6}~a^{9/6}~b^{6/6}~c^{3/6}}\] \[\Huge{27^{1/6}~a^{1~~1/2}~b^{1}~c^{1/2}}\] that mixed numeral, the reason why we want it is it tells us how to split up that "a": a^1 and a^(1/2) it also might be good to know that 27 = 3^6, giving us 3^(6/6) \[\Huge{3ab~a^{1/2}~c^{1/2}}\]

OpenStudy (amistre64):

since the denominator of an exponent is the radical number ... we can rewrite those 1/2s as sqrts if need be

OpenStudy (amistre64):

pfft, i got my 27^(1/6) bad didnt i ...

OpenStudy (amistre64):

27 = 3^3; giving us 3^(3/6), or 3^(1/2)

OpenStudy (anonymous):

Okay, so the final answer would be 3^{1/2} ?

OpenStudy (anonymous):

Or do we need to clarify more?

OpenStudy (amistre64):

3^(1/2) is only the one factor that I messed up on along the way .... the rest of them are still a part of the solution :/

OpenStudy (amistre64):

your setup simplifies to: \[\Huge{ab~3^{1/2}~a^{1/2}~c^{1/2}}\] you might want to tuck those 1/2 parts under a radical again to complete the process

OpenStudy (anonymous):

What do you mean 'tuck them under a radical again'?

OpenStudy (amistre64):

werent you here when we went over the property:\[\Huge \sqrt[k]{a^n}=a^{(n/k)}\] we can convert a radical to a fractional exponent; and we can also convert a fractional exponent back into a radical. A radical has something underneath it, tucked inside of it, placed within it, crammed betwixt it .... etc.

OpenStudy (anonymous):

Oh okay, sorry. Blonde moment xD. Thanks so much :)!

OpenStudy (amistre64):

youre welcome. if you want to post your final result, i can chk it for you.

OpenStudy (anonymous):

Well, that was the correct answer, did not need simplify further, thanks again so much.